General Electronics produces electronic communications equipments. In 2014 the company had Tk. 200,000 earnings before interest and taxes. On January 1, 2014 the company borrowed Tk. 400,000 at a rate of interest of 10 percent. The company had no previous debt and its tax rate is 40 percent.

1.What was the degree of financial leverage prior to 2014?

2.What was the 2014 degree of financial leverage using the actual earnings figures?

3.In 2014 what percentage change in after tax earnings per share would result from a 50 percent increase in EBIT?

4.In 2014 what percentage increase in EBIT would bring about a 20 percent increase in EPS?

5.In 2014 the company sold 20,000 items at Tk. 60 per item. The fixed operating cost was Tk. 3600,000, and the total variable operating cost was Tk. 400,000. What was the degree of combined leverage using the actual 2014 earnings figures?

6.Compute DOL and check the relation DCL = DOL X DFL using the data and results of ii and v.

P&G has EBIT of $67,500. Interest costs are $22,500 and the firm has 15,000 shares of common stock outstanding. Assume a 40% tax rate.

7.Calculate the financial leverage for the firm.

8.If the firm also has 1,000 shares of preferred stock paying a $6 annual dividend per share, what is the DFL?

Play-More Toys produces beach balls, selling 400,000 balls a year. Each ball produced has a variable operating cost of Tk. 7.5 and sells for Tk. 10. Fixed operating costs are Tk. 280,000. The firm has annual interest charges of Tk. 60,000, preferred dividends of Tk. 20,000 and a 40% tax rate.

9.Calculate the operating breakeven point in unit.

10.Calculate DOL, DFL and DTL
Answer #1

The formula for Degree of Financial Leverage is EBIT/EBIT-Interest Expense

For first question prior to 2014, the interest expense is zero. So plugging in the values in the formula we get 1 as the answer.

For second question EBIT is 200,000 while interest expense is 10% of 400,000 = Tk 40,000

Therefore, Degree of Financial Leverage is (200,000)/(200,000-40,000) = 1.25

For third question, The EBIT increases by 50%, (the growth in EPS in % will be equal as the growth in Net After Tax Return or PAT)

the earlier EBIT was 200,000 after interest it would be 160,000 so after tax it would be 160,000 * (1-0.4) = 96,000

Now as the EBIT grew by 50% the EBIT is 300,000, after interest it would be 260,000  and thus after tax would be 260,000 * (1-0.4) = 156,000 Thus the EPS has grew 62.5%  from 96,000 to 156,000

For Fourth Question the EPS or Net income will grow to 115,200 thus now adding back 40% tax = 115200/0.6 = 192,000

Now adding back interest 40,000 to EBT = 192000 + 40000 = 232,000 which is 16% higher than earlier EBIT of 200,000